# Ashish Saxena 11    Tanvir Kaur 11    Kaushik Mondal 11

### v2

@@ 3.1 Algorithm for general graph
Suppose $n+1$ agents are co-located at a node say $v$ of the graph (rooted configuration). Initially, for each agent $a_{i}$, $a_{i}.grp\_label=10$. Therefore, all agents are a part of $G_{1}$. According to our model, an agent cannot sense the presence of a missing edge unless it attempts to move through it. To tackle this, we take two consecutive rounds to perform one edge traversal. In every odd round, agents move, and in the next even round, it realizes the movement is successful or unsuccessful. Precisely, say an agent $a_{i}$ attempts to move in an odd round $r$. If its movement is successful, it updates its parameter $a_{i}.success$ to $True$ in the even round $r+1$. Otherwise, it updates $a_{i}.success=False$ in the even round $r+1$. Now we proceed with a detailed description of our algorithm.
Algorithm for unsettled agent $\bm{a_{i}}$ of $\bm{G_{1}}$ at odd round: In any odd round, let agent $a_{i}\in G_{1}$ be at node $u$ and $a_{i}.settled=0$. It follows the following steps.
  • $\bm{a_{i}.divide=0}$: Suppose, agent $a_{i}$ reaches a node $w$ and finds no agents $a_{j}$ with $a_{j}.settled=1$ (i.e., no settled agents at node $w$). In this case, it follows the following steps. If agent $a_{i}$ is the minimum ID at node $w$, then it settles at $w$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$. Further, it updates $a_{i}^{w}(G_{1})=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$. Else if agent $a_{i}$ is not the minimum ID, then it moves according to the DFS traversal algorithm. If agent $a_{i}$ reaches a node $v$ and finds an agent with $a_{j}$ with $a_{j}.settled=1$, then it $a_{i}$ does not settle at node $v$ and moves according to the DFS traversal algorithm. In this way, the agent $a_{i}\in G_{1}$ continues the DFS traversal of the graph unless $a_{i}$ encounters a missing edge for the first time. In particular, at some odd round $r$, agent $a_{i}$ moves via an edge $e$, and the movement is not successful at the end of round $r$. This is understood by agent $a_{i}$ in the round $r+1$. In round $r+2$, agent $a_{i}$ updates $divide$ by 1 and divides into two groups using the $Group\_divide$ procedure as follows.
@@
  • $\bm{a_{i}.divide=1}$: Suppose, agent $a_{i}\in G_{1}$ reaches a node $w$ and finds no agents $a_{j}$ with $a_{j}.settled=1$ (i.e., no settled agents at node $w$). In this case, it follows the following steps. If agent $a_{i}$ is the minimum ID at node $w$, then it settles at $w$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$. Further, it updates $a_{i}^{w}(G_{1})=(a_{i}.prt\_in,\,a_{i}.grp\_label,\,a_{i}.dfs\_label)$. If there is any unsettled agent $a_{j}$ from $G_{2}$ at node $w$, then $a_{i}$ updates $a_{i}^{u}(G_{2})$ to $(a_{j}.prt\_in,a_{j}.grp\_label,a_{j}.dfs\_label)$. If $a_{i}\in G_{1}$ reaches a node $w$ and find agent $a_{j}$ with $a_{j}.settled=1$, then it does not settle at the node $v$ and it can understand whether node $w$ is the new node according to the current DFS or previously visited node by comparing $label$, $dfs\_label$ components of $a_{j}^{w}(G_{1})$ with $a_{i}.grp\_label$, $a_{i}.dfs\_label$ respectively. If both matches then node $w$ is visited previously in the current DFS traversal of agent $a_{i}$. Otherwise, if any of those two components do not match, then agent $a_{i}\in G_{1}$ treats $w$ as an unvisited node w.r.t. the current DFS. Round $r+2$ onwards, $a_{i}$ updates its $count_{1}$ parameter as follows. At odd round $r_{1}\geq r+2$, if it finds $a_{i}.success=True$, i.e. it realized a successful movement in the even round $r_{1}-1$, agent $a_{i}$ updates $a_{i}.count_{1}=0$; else if $a_{i}.success=False$, agent $a_{i}$ updates $a_{i}.count_{1}=a_{i}.count_{1}+1$. If at some odd round, $a_{i}$ finds $a_{i}.count_{1}=16n^{2}$, then unsettled agents at node $u$ divide into two groups again using aforementioned $Group\_divide$ procedure. That is, $G_{1}$ divides into two groups only after getting stuck at a node for a particular port for $16n^{2}$ many odd rounds.
  • (1)
    $\bm{a_{i}.state=explore}$:
@@
      $\bm{a_{i}.success=True}$: It updates $a_{i}.count_{2}=0$ and does the following.
@@
        Node $\bm{u}$ has no agent with $\bm{setlled=1}$: If $a_{i}$ is the minimum ID at node $u$, it settles at node $u$ and updates $a_{i}.settled=1$, $a_{i}.grp\_label=\bot$ and $a_{i}^{u}(G_{2})$ to $(a_{i}.prt\_in,a_{i}.grp\_label,a_{i}.dfs\_label)$. If there is any unsettled agent $a_{j}$ from $G_{1}$ at node $u$, then $a_{i}$ updates $a_{i}^{u}(G_{1})$ to $(a_{j}.prt\_in,a_{j}.grp\_\\ label,a_{j}.dfs\_label)$. If $a_{i}$ is not the minimum ID, then it updates $a_{i}.prt\_out=(a_{i}.prt\_in+1)$ mod $\delta_{u}$. If $a_{i}.prt\_out\neq a_{i}.prt\_in$, then it moves through $a_{i}.prt\_out$. Else, if $a_{i}.prt\_out=a_{i}.prt\_in$, then it sets $a_{i}.state=backtrack$ and move through $a_{i}.prt\_out$.

### v3

@@ 6 Impossibility and lower bound for $\ell$-bounded 1-interval connected graphs
      $=S_{P_{r}}(w_{4})+x_{3}-y_{4}+(y_{5}-x_{4})-(y_{5}-x_{4})$
      $=S_{P_{r}}(w_{4})+(y_{5}+x_{3})-(y_{4}+x_{4})+(x_{4}-y_{5})$
      $=p+(x_{4}-y_{5})-c_{2}\;\;$
      $\text{ as }\;S_{P_{r}}(w_{4})+(y_{5}+x_{3})-(y_{4}+x_{4})=E_{P_{r}}(w_{4})=p-c_{2}$
      $\therefore E_{P_{r}^{\prime}}(w_{4})$ $\geq p+2,\;\text{ as }\;x_{4}-y_{5}\geq 2+c_{2}$
    • (b)
      If $y_{4}^{\prime}\geq y_{4}+1$,
      $E_{P_{r}^{\prime}}(w_{3})$ $=S_{P_{r}^{\prime}}(w_{3})+(y_{4}^{\prime}+x_{2})-(y_{3}+x_{3})$
@@
      $E_{P_{r}^{\prime}}(w_{3})$ $\geq S_{P_{r}}(w_{3})+(y_{4}+x_{2})-(y_{3}+x_{3})+1,\;\text{ as }\;y_{4}^{\prime}\geq y_{4}+1$
      $=p+2,\;\text{ as }\;S_{P_{r}}(w_{3})+(y_{4}+x_{2})-(y_{3}+x_{3})=E_{P_{r}}(w_{3})=p+1$
      $\therefore E_{P_{r}^{\prime}}(w_{3})$ $\geq p+2$
  • 3.
    Case 3: In this case, $x_{4}-y_{5}\geq 2+c_{1}$. We divide this case into the following sub-cases.
@@
      If $y_{4}^{\prime}\leq y_{4}+c_{1}+1$,
      $E_{P_{r}^{\prime}}(w_{4})$ $=S_{P_{r}^{\prime}}(w_{4})+x_{3}-y_{4}^{\prime}$
[…]
      $E_{P_{r}^{\prime}}(w_{4})$ $\geq S_{P_{r}}(w_{4})+x_{3}-y_{4}-c_{1}-1\;\text{ as }\;y_{4}^{\prime}\leq y_{4}+c_{1}+1$
      $=S_{P_{r}}(w_{4})+x_{3}-y_{4}+(y_{5}-x_{4})-(y_{5}-x_{4})-c_{1}-1$
      $=S_{P_{r}}(w_{4})+(y_{5}+x_{3})-(y_{4}+x_{4})+(x_{4}-y_{5})-c_{1}-1$
      $=p+(x_{4}-y_{5})-c_{1}\;\;\text{ as }\;S_{P_{r}}(w_{4})+(y_{5}+x_{3})-(y_{4}+x_{4})=S_{P_{r}}(w_{4})=p+1$
      $\therefore E_{P_{r}^{\prime}}(w_{4})$ $\geq p+2\;\text{ as }\;x_{4}-y_{5}\geq 2+c_{1}$
    • (b)
      If $y_{4}^{\prime}\geq y_{4}+c_{1}+2$,
      $E_{P_{r}^{\prime}}(w_{3})$ $=S_{P_{r}^{\prime}}(w_{3})+(y_{4}^{\prime}+x_{2})-(y_{3}+x_{3})$
[…]
      $E_{P_{r}^{\prime}}(w_{3})$ $\geq S_{P_{r}}(w_{3})+(y_{4}+x_{2})-(y_{3}+x_{3})+c_{1}+2,\;\text{ as }\;y_{4}^{\prime}\geq y_{4}+c_{1}+2$
      $=p+2,\;\text{ as }\;S_{P_{r}}(w_{3})+(y_{4}+x_{2})-(y_{3}+x_{3})=E_{P_{r}}(w_{3})=p-c_{1}$
      $\therefore E_{P_{r}^{\prime}}(w_{3})$ $\geq p+2$
  • 4.
    Case 4: In this case, $x_{4}-y_{5}\geq 1$. We divide this case into the following sub-cases.
@@
      If $y_{4}^{\prime}\leq y_{4}$,
      $E_{P_{r}^{\prime}}(w_{4})$ $=S_{P_{r}^{\prime}}(w_{4})+x_{3}-y_{4}^{\prime}$
[…]
      $=S_{P_{r}}(w_{4})+x_{3}-y_{4}+(y_{5}-x_{4})-(y_{5}-x_{4})$
      $=S_{P_{r}}(w_{4})+(y_{5}+x_{3})-(y_{4}+x_{4})+(x_{4}-y_{5})$
      $E_{P_{r}^{\prime}}(w_{4})$ $\geq p+1+(x_{4}-y_{5}),\;\;\text{ as }\;S_{P_{r}}(w_{4})+(y_{5}+x_{3})-(y_{4}+x_{4})=E_{P_{r}}(w_{4})=p+1$
      $\therefore\;E_{P_{r}^{\prime}}(w_{4})$ $\geq p+2,\;\text{ as }\;x_{4}-y_{5}\geq 1$